If 20 g steam initially at 1000C is added to 60 g of ice initially at 00C, then find the final equilibrium temperature of the mixture.


1
Expert's answer
2021-08-16T08:36:11-0400

The heat energy lost by the steam is equal to the heat energy gained by the ice.

The expression for the loss or gain in heat energy during a phase change is

"Q=mL_f"

Here, m is mass, Lf is latent heat of fusion.

The expression for the loss or gain in heat energy during a phase change is

"Q=mL_v"

Here, m is mass, Lv is latent heat of vaporization.

The heat energy required to change the phase of ice at constant temperature 0 °C is

"Q_1=m_{ice}L_f"

The heat energy required to raise the temperature of water from 0 °C to equilibrium temperature te is

"Q_2=m_{ice}c_w(t_e – 0 \;°C)"

Here, mice is mass of ice and cw is specific heat of water.

The heat energy required to change the phase of steam at constant temperature 100 °C is

"Q_3 = m_{steam}L_v"

Here, msteam is mass of steam and Lv is latent heat of vaporization.

The heat energy required to lower the temperature of water from 100 °C to equilibrium temperature te is

"Q_4=m_{steam}c_w(100 \;°C -t_e)"

The heat energy lost by the steam is equal to the heat energy gained by the ice. Hence

Q_1+Q_2=Q_3+Q_4

Substitute the values of Q1, Q2, Q3 and Q4 in the above equation to solve for te.

"m_{ice}L_f + m_{ice}c_w(t_e – 0 \;°C) = m_{steam}L_v + m_{steam}c_w(100 \;°C -t_e) \\

m_{ice}c_wt_e+m_{steam}c_wt_e= m_{steam}L_v + m_{steam}c_w100\;°C -m_{ice}L_f \\

t_e(m_{ice}c_w + m_{steam}c_w) = m_{steam}L_v + m_{steam}c_w100\;°C -m_{ice}L_f \\

t_e = \frac{m_{steam}L_v + m_{steam}c_w100\;°C -m_{ice}L_f}{m_{ice}c_w + m_{steam}c_w} \\

m_{ice} = 60 \;g \\

m_{steam} =20 \;g \\

L_v=540 \;cal/g \\

c_w = 1.0 \;cal/g \cdot °C \\

L_f = 80 \;cal/g \\

t_e = \frac{20 \times 540 + 20 \times 1.0 \times 100 -60 \times 80 }{60 \times 1.0 + 20 \times 1.0} \\

= \frac{10800 + 2000 -4800}{80} \\

= 100 \;°C"

Therefore, the equilibrium temperature of the system is 100 °C.